answer all

R; Problem 2) What are the values of Rin, Rout, Ay, and A, for the amplifier given if R = 5000, R3 = 1MA, RE = 120k2, Rc =
Answer #1

- at small signal - 9m = parameters Jeg G.T. a point (Icq VLEO) = (0.109mA, 12-20) IC - BIB IB = IC IB = 0-100 mA 100 IB 1-09

- small signal model Shie RA I l Iolithfe) &RE current gain AI I AI = -hfe to AI = -hee AJ - -100 AA at Alp loop sy Anput resPihte = B=100] hie = 23.64 kr hi - (23.648 +(1+100) 1208) К; - 12-144 x16 л Ri = 12.144 mm voltage gain Av = we are the = hfeAu rotal = Av. (3) A vende z u einem Apply voltage divinon at vi - S X / Re RA TRB v = RB RothB Avrobal - Av. (4) {here to ge